# Internal Consistency and Self-Feedback in Large Language Models: A Survey

### v1

@@ Abstract
Large language models (LLMs) are expected to respond accurately but often exhibit deficient reasoning or generate hallucinatory content. To address these, studies prefixed with “Self-” such as Self-Consistency, Self-Improve, and Self-Refine have been initiated. They share a commonality: involving LLMs evaluating and updating itself to mitigate the issues. Nonetheless, these efforts lack a unified perspective on summarization, as existing surveys predominantly focus on categorization without examining the motivations behind these works.
In this paper, we summarize a theoretical framework, termed Internal Consistency, which offers unified explanations for phenomena such as the lack of reasoning and the presence of hallucinations. Internal Consistency assesses the coherence among LLMs’ latent layer, decoding layer, and response layer based on sampling methodologies. Expanding upon the Internal Consistency framework, we introduce a streamlined yet effective theoretical framework capable of mining Internal Consistency, named Self-Feedback. The Self-Feedback framework consists of two modules: Self-Evaluation and Self-Update. The former captures Internal Consistency signals, while the latter leverages the signals to enhance either the model’s response or the model itself. This framework has been employed in numerous studies.
We systematically classify these studies by tasks and lines of work; summarize relevant evaluation methods and benchmarks; and delve into the concern, “Does Self-Feedback Really Work?” We propose several critical viewpoints, including the “Hourglass Evolution of Internal Consistency”, “Consistency Is (Almost) Correctness” hypothesis, and “The Paradox of Latent and Explicit Reasoning”. Furthermore, we outline promising directions for future research. We have open-sourced the experimental code, reference list, and statistical data, available at https://github.com/IAAR-Shanghai/ICSFSurvey.
@@ I Introduction
Large language models (LLMs), including ChatGPT [^1] https://openai.com/index/chatgpt/, have made significant advancements in the field of natural language processing (NLP), demonstrating capabilities close to basic human intelligence. These include features like performing basic reasoning and learning from examples [1].
@@
Figure 1: GPT-4o provides different answers to the same question. The complete responses can be found in Appendix B.
@@ II-A Formulation
Consistency is a critical term in logic, referring to a system where no two statements contradict each other [31]. However, systems like those of language models typically exhibit inconsistencies. For instance, as shown in Fig. 1, even GPT-4o can’t guarantee consistent responses to the same question. To better define the internal consistency, we utilize a sampling-based approach to model expressions in LLMs [25] [^4] There are various methods to model the Internal Consistency, and the sampling-based perspective is a common one derived from the existing works. Important symbols discussed in this paper can be referred to in Appendix A..
@@ III-B Taxonomy
In terms of specific manifestations, low Internal Consistency includes two major issues: weak reasoning ability in QA tasks and hallucinations in open-ended generation. Therefore, in Section V, we summarize three lines of work for Reasoning Elevation. And in Section VI, we summarize four lines of work for Hallucination Alleviation. Table III summarizes these seven lines of work. Additionally, some tasks do not aim to improve consistency but still utilize the Self-Feedback framework. For comprehensiveness, we also briefly summarize these other tasks in Section VII.
TABLE III: Different Lines of Work in Reasoning Elevation and Hallucination Alleviation
@@ IV-B Confidence Estimation
Prompt for Self-Evaluation Question: Who was the first president of the United States?
@@ V Task: Reasoning Elevation
Reasoning Elevation refers to enhancing the logical reasoning capabilities of language models during response generation to improve their internal consistency. The primary feature of this line of work is the use of benchmarks in the form of QA tasks. We have identified three significant lines of work, as shown in the upper part of Table III.
@@ VI Task: Hallucination Alleviation
Reasoning elevation typically targets QA tasks, while hallucination alleviation is generally aimed at open-ended generation tasks such as story writing and code generation, emphasizing goals like fact enhancement, error reduction, and faithfulness enhancement. This section provides a survey of these directions. We have categorized four significant lines of work, as shown in the lower half of Table III.
@@ VI-A Refining the Response Iteratively
This line of work is similar to Refining with Responses listed in Section V-B. The difference lies in the focus: Refining with Responses primarily targets QA tasks, where various methods often only need to consider intermediate reasoning steps and the correctness of the final answer. This can be achieved by sampling multiple responses and synthesizing or selecting the better ones. In contrast, hallucination alleviation primarily deals with open-ended tasks such as story generation and code generation. When sampling multiple responses, it is still necessary to meticulously check each response for errors. Therefore, to alleviate hallucinations, it is crucial to iteratively refine and polish a response, eliminating errors. Their comparison is shown in Table IV.
TABLE IV: Refining with Responses V.S. Refining the Response Iteratively
@@ VII-A Preference Learning
There are three main ways to obtain preference feedback. The first is through human feedback, as seen in works like OASST [113] and BeaverTails [114], which include human-annotated data from multiple annotators. The second method involves feedback generated by models [115, 116, 117], offering lower annotation costs and faster iterative feedback efficiency compared to human feedback. The third type of feedback is derived from inductive bias. For instance, the SHP dataset mentioned in [118] considers upvotes and downvotes to differentiate between good and bad responses. Another example is ALMoST [119], which proposes some prior rules, such as larger parameter models generally performing better, and more prompt context examples yielding higher quality responses. These rules are used to rank response quality across different models as preference feedback.
@@ VIII-A Meta Evaluation
We summarize five meta evaluation methods. These methods can generally be categorized into metric-based and benchmark-based paradigms. The former constructs mathematical formulas or metrics to directly calculate the performance of a particular aspect; the latter uses QA datasets to empirically measure performance. We summarize benchmark-based meta evaluations in Table V.
TABLE V: Meta Evaluation Benchmarks
@@ VIII-B Common Evaluation
In most works utilizing the Self-Feedback framework, they do not conduct meta evaluations in experiments but rather assess the model’s ability to solve real-world problems, such as solving math problems, code generation, and summarizing articles. Here, we classify and summarize the popular benchmarks based on the focus of the evaluation tasks [^17] We may see the same benchmark categorized differently in different sources, highlighting the difficulty in achieving fully orthogonal classifications. For example, a dataset evaluating linguistic understanding ability may also involve reasoning tasks. Here, we classify each dataset based on its primary focus., as shown in Table VI. There are two types of reasoning: Knowledge reasoning focuses on the LLM’s ability to solve problems using its parameterized knowledge, while Logic reasoning focuses on complex logical reasoning based on prompts. Linguistic understanding focuses on the LLM’s analysis of language meaning. Code generating and Math Solving respectively test the LLM’s ability to write code and solve math problems.
TABLE VI: Common Evaluation Benchmarks
@@
Knowledge reasoning MMLU [15, 139] UCB
@@ IX-A Conflicting Viewpoints
    Does Self-Feedback improve Internal Consistency?
@@
    Does Internal Consistency mean correctness?
IX-B Does Self-Feedback Improve Internal Consistency?
For this question, the different lines of work proposed in this paper actually provide affirmative answers from various perspectives. For example,
  • Self-Consistency [2] (Reasoning Topologically) improves response consistency by finding the majority vote from the sampling set.
  • Self-Improve [46] (Refining with Responses) also uses majority voting to find better responses from the sampling set and then fine-tunes the model itself, improving consistency from the parameter level.
  • Multi-Agent Debate [73] (Multi-Agent Collaboration) involves multiple models participating in answering questions simultaneously, which is essentially a generalized way to obtain consistent responses.
  • Self-Refine [8] (Refining the Response Iteratively) iteratively optimizes its responses, which is an implicit way to achieve high-consistency responses.
  • Self-Contradict [21] (Mitigating Hallucination while Generating) eliminates self-contradictory content in its responses, naturally improving consistency.
  • DoLa [55] (Decoding Truthfully) compares the probability distributions across different layers of the model’s latent layers and reduces their discrepancies through comparison, thereby improving the model’s consistency.
  • ITI [5] (Activating Truthfulness) implicitly optimizes the model’s Internal Consistency by identifying attention heads that prefer factual information.
IX-C Does Internal Consistency Mean Correctness?
@@ X-E The Comprehensive Evaluation
XI Limitations
Given the terminological confusion and overlapping lines of work in this field, the sections of this paper are not necessarily orthogonal. Consequently, a single work may fall into multiple categories, indicating that the method employs ideas from different lines of work. The writing design is more inspirational, helping researchers to answer: What is the position of their work in a grand system? What are its advantages and disadvantages? What should be the next step in designing new and better methods?
XII Conclusion
@@
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grants No. 62072463, 71531012), the National Social Science Foundation of China (Grants No. 18ZDA309), the Research Seed Funds of the School of Interdisciplinary Studies at Renmin University of China, and the Opening Project of the State Key Laboratory of Digital Publishing Technology of the Founder Group.
Appendix A Notations
TABLE VII: Common Notations
Symbol Description
$\bm{x}$ Query
$\mathcal{M},\mathcal{N}$ LLMs
$O_{e}(\mathcal{M},\bm{x})$ Sampling distribution
$\mathcal{Y}$ Sampling set
$\bm{y}_{i}$ The $i$-th element in the sampling set
$\bm{y}_{0:i}$ Elements from 00 to $i$ in the sampling set
$y^{t}$ The $t$-th token in text $\bm{y}$
$f$ Consistency Signal of Self-Feedback
$P(\bm{y}|\bm{x};\theta)$ Probability that model parameterized by $\theta$ outputs $\bm{y}$ given input $\bm{x}$
Appendix B The Five Responses of GPT-4o
Fig. 10,  11,  12,  13, and 14 show five different responses obtained when querying GPT-4o: How many full stops (periods) are there: “.!..!..!”. The experiments were conducted on June 14, 2024, accessed via the official website [^20] https://chat.openai.com/.
GPT-4o Response 1 To determine the number of full stops (periods) in the string ”.!..!..!”, we need to count the occurrences of the character ’.’.
Let’s break down the string:
1. ”.!”: 1 full stop
2. ”..!”: 2 full stops
3. ”.!”: 1 full stop
Summing these up:
1 + 2 + 1 = 4
Therefore, there are 4 full stops in the string ”.!..!..!”.
Figure 10: GPT-4o Response 1
GPT-4o Response 2 The sequence ”.!..!..!” contains three instances of ”.” (full stops or periods). Each segment ”.!” has one full stop, and there are three such segments.
Figure 11: GPT-4o Response 2
GPT-4o Response 3 The string ”.!..!..!” contains a total of **3** full stops (periods).
Figure 12: GPT-4o Response 3
GPT-4o Response 4 The string ”.!..!..!” contains full stops (periods). To determine the number of full stops in this string, we simply count the occurrences of the period character (’.’). Let’s analyze the string:
- The first character is a period.
- The second character is an exclamation mark.
- The third character is a period.
- The fourth character is an exclamation mark.
- The fifth character is a period.
- The sixth character is an exclamation mark.
- The seventh character is a period.
- The eighth character is an exclamation mark.
Counting the periods:
1. The first period is at position 1.
2. The second period is at position 3.
3. The third period is at position 5.
4. The fourth period is at position 7.
Therefore, there are 4 full stops (periods) in the string ”.!..!..!”.
Figure 13: GPT-4o Response 4
GPT-4o Response 5 The sequence ”.!..!..!” contains the following characters:
1. Period (.)
2. Exclamation mark (!)
3. Period (.)
4. Exclamation mark (!)
5. Period (.)
6. Exclamation mark (!)
To count the full stops (periods):
1. The first character is a period.
2. The third character is a period.
3. The fifth character is a period.
There are a total of **three** periods in the sequence ”.!..!..!”.
Figure 14: GPT-4o Response 5
Appendix C Experiment Details of Three Types of Consistency
The setups and results of the comparative experiments on three different types of consistency are shown in Tables VIII,  IX, and  X. Here, Fix $\text{attn}^{16}_{15}$ refers to keeping the 16th attention head in the 15th layer unchanged, while zero out $\text{attn}^{i\neq 16}_{15}$ denotes zeroing out all attention heads in the 15th layer except the 16th one. The source code for this experiment is available in our open-source GitHub repository.
TABLE VIII: Latent Consistency
Setting Selected Token
Fix $\text{attn}^{0}_{0}$; Zero out $\text{attn}^{i\neq 0}_{0}$ 0
Fix $\text{attn}^{16}_{0}$; Zero out $\text{attn}^{i\neq 16}_{0}$ 0
Fix $\text{attn}^{0}_{15}$; Zero out $\text{attn}^{i\neq 0}_{15}$ 5
Fix $\text{attn}^{16}_{15}$; Zero out $\text{attn}^{i\neq 16}_{15}$ 4
Fix $\text{attn}^{0}_{30}$; Zero out $\text{attn}^{i\neq 0}_{30}$ 4
Fix $\text{attn}^{16}_{30}$; Zero out $\text{attn}^{i\neq 16}_{30}$ 4
TABLE IX: Decoding Consistency
Decoding Strategy Selected Token
Greedy Decoding 4
Beam Search Decoding ($\text{beam}=5$) 4
Sampling Decoding 3
Top-k Sampling Decoding ($k=50$) 4
Top-p Sampling Decoding ($p=0.95$) 4
TABLE X: Response Consistency
Model Response
RESPONSE 1: There is a period after the first dot, then another period after the second dot, and so on. So, there are 5 periods in total.
RESPONSE 2: There is one full stop after the first ”!”, then another after the second ”!”, and another after the third ”!”. So, there are 3 full stops in total.
RESPONSE 3: There is one full stop after the first ”!”, then another after the second ”!”, and so on. So, there are 3 full stops.
RESPONSE 4: There is one period at the end of the first part, then another after the second part, and finally one more after the third part. So, there are 3 periods in total.
RESPONSE 5: There is 1 period, then another one, and another one… So, there are 3 full stops!

### v2

@@ Abstract
Large language models (LLMs) often exhibit deficient reasoning or generate hallucinations. To address these, studies prefixed with “Self-” such as Self-Consistency, Self-Improve, and Self-Refine have been initiated. They share a commonality: involving LLMs evaluating and updating themselves. Nonetheless, these efforts lack a unified perspective on summarization, as existing surveys predominantly focus on categorization.
In this paper, we summarize a theoretical framework, Internal Consistency, offering explanations for reasoning deficiencies and hallucinations. Internal Consistency refers to the consistency in expressions among LLMs’ latent, decoding, or response layers based on sampling methodologies. Then, we introduce another effective theoretical framework capable of mining Internal Consistency, named Self-Feedback. This framework consists of two modules: Self-Evaluation and Self-Update. The former captures Internal Consistency Signals, while the latter leverages the signals to enhance either the model’s response or the model itself. This framework has been employed in numerous studies.
We systematically classify these studies by tasks and lines of work; summarize relevant evaluation methods and benchmarks; and delve into the concern, “Does Self-Feedback Really Work?” We also propose several critical viewpoints, including the “Hourglass Evolution of Internal Consistency”, “Consistency Is (Almost) Correctness” hypothesis, and “The Paradox of Latent and Explicit Reasoning”. The relevant resources are open-sourced at https://github.com/IAAR-Shanghai/ICSFSurvey.
@@ I Introduction
Large language models (LLMs) have made significant advancements in the field of natural language processing (NLP), demonstrating capabilities close to basic human intelligence. These include features like performing basic reasoning and learning from examples [1].
@@
Figure 1: GPT-4o provides different answers to the same question. The complete responses can be found in our GitHub open-source repository.
@@ II-A Formulation
Consistency is a critical term in logic, referring to a system where no two statements contradict each other [31]. However, systems like those of language models typically exhibit inconsistencies. For instance, as shown in Fig. 1, even GPT-4o can’t guarantee consistent responses to the same question. To better define the internal consistency, we utilize a sampling-based approach to model expressions in LLMs [25]. In addition, Table III provides explanations of some notations frequently used in this paper.
TABLE III: Common Notations
Symbol Description
$\bm{x}$ Query
$\mathcal{M},\mathcal{N}$ LLMs
$O_{e}(\mathcal{M},\bm{x})$ Sampling distribution
$\mathcal{Y}$ Sampling set
$\bm{y}_{i}$ The $i$-th element in the sampling set
$\bm{y}_{0:i}$ Elements from 00 to $i$ in the sampling set
$y^{t}$ The $t$-th token in text $\bm{y}$
$f$ Consistency Signal of Self-Feedback
$P(\bm{y}|\bm{x};\theta)$ Language model parameterized by $\theta$
@@ III-B Taxonomy
In terms of specific manifestations, low Internal Consistency includes two major issues: weak reasoning ability in QA tasks and hallucinations in open-ended generation. Therefore, in Section V, we summarize three lines of work for Reasoning Elevation. And in Section VI, we summarize four lines of work for Hallucination Alleviation. Table IV summarizes these seven lines of work. Additionally, some tasks do not aim to improve consistency but still utilize the Self-Feedback framework. For comprehensiveness, we also briefly summarize these other tasks in Section VII.
TABLE IV: Different Lines of Work in Reasoning Elevation and Hallucination Alleviation
@@ IV-B Confidence Estimation
Question: Who was the first president of the United States?
@@ V Task: Reasoning Elevation
Reasoning Elevation refers to enhancing the logical reasoning capabilities of language models during response generation to improve their internal consistency. The primary feature of this line of work is the use of benchmarks in the form of QA tasks. We have identified three significant lines of work, as shown in the upper part of Table IV.
@@ VI Task: Hallucination Alleviation
Reasoning elevation typically targets QA tasks, while hallucination alleviation is generally aimed at open-ended generation tasks such as story writing and code generation, emphasizing goals like fact enhancement, error reduction, and faithfulness enhancement. This section provides a survey of these directions. We have categorized four significant lines of work, as shown in the lower half of Table IV.
@@ VI-A Refining the Response Iteratively
This line of work is similar to Refining with Responses listed in Section V-B. The difference lies in the focus: Refining with Responses primarily targets QA tasks, where various methods often only need to consider intermediate reasoning steps and the correctness of the final answer. This can be achieved by sampling multiple responses and synthesizing or selecting the better ones. In contrast, hallucination alleviation primarily deals with open-ended tasks such as story generation and code generation. When sampling multiple responses, it is still necessary to meticulously check each response for errors. Therefore, to alleviate hallucinations, it is crucial to iteratively refine and polish a response, eliminating errors. Their comparison is shown in Table V.
TABLE V: Refining with Responses V.S. Refining the Response Iteratively
@@ VII-A Preference Learning
There are three main ways to obtain preference feedback. The first is through human feedback, as seen in works like OASST [113] and BeaverTails [114], which include human-annotated data from multiple annotators. The second method involves feedback generated by models [115, 116], offering lower annotation costs and faster iterative feedback efficiency compared to human feedback. The third type of feedback is derived from inductive bias. For instance, the SHP dataset mentioned in [117] considers upvotes and downvotes to differentiate between good and bad responses. Another example is ALMoST [118], which proposes some prior rules, such as larger parameter models generally performing better, and more prompt context examples yielding higher quality responses. These rules are used to rank response quality across different models as preference feedback.
@@ VIII-A Meta Evaluation
We summarize five meta evaluation methods. These methods can generally be categorized into metric-based and benchmark-based paradigms. The former constructs mathematical formulas or metrics to directly calculate the performance of a particular aspect; the latter uses QA datasets to empirically measure performance. We summarize benchmark-based meta evaluations in Table VI.
TABLE VI: Meta Evaluation Benchmarks
@@ VIII-B Common Evaluation
In most works utilizing the Self-Feedback framework, they do not conduct meta evaluations in experiments but rather assess the model’s ability to solve real-world problems, such as solving math problems, code generation, and summarizing articles. Here, we classify and summarize the popular benchmarks based on the focus of the evaluation tasks [^15] We may see the same benchmark categorized differently in different sources, highlighting the difficulty in achieving fully orthogonal classifications. For example, a dataset evaluating linguistic understanding ability may also involve reasoning tasks. Here, we classify each dataset based on its primary focus., as shown in Table VII. There are two types of reasoning: Knowledge reasoning focuses on the LLM’s ability to solve problems using its parameterized knowledge, while Logic reasoning focuses on complex logical reasoning based on prompts. Linguistic understanding focuses on the LLM’s analysis of language meaning. Code generating and Math Solving respectively test the LLM’s ability to write code and solve math problems.
TABLE VII: Common Evaluation Benchmarks
@@
Knowledge reasoning MMLU [15] UCB
@@ IX-A Conflicting Viewpoints
    Does Self-Feedback improve Internal Consistency? The answer is yes. As demonstrated in our survey, different lines of research offer affirmative evidence from various perspectives.
@@
    Does Internal Consistency mean correctness? We cannot directly conclude this. We will delve deeper into this question in the following section.
IX-B Does Internal Consistency Mean Correctness?
@@ X-E The Comprehensive Evaluation
XI Conclusion